\begin{thebibliography}{10}

%\cite{Bakeyev:2001wb}%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\bibitem{Bakeyev:2001wb}
T.~Bakeyev,
%``A new way to deal with fermions in Monte Carlo simulations,''
hep-lat/0105008, to appear in Phys.\ Lett.\ B.
%%CITATION = HEP-LAT 0105008;%%

%\cite{HMC}%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\bibitem{HMC}
S.~Duane et. al.
%A.~D.~Kennedy, B.~J.~Pendleton and D.~Roweth,
%``Hybrid Monte Carlo,''
Phys.\ Lett.\ B {\bf 195} (1987) 216.
%%CITATION = PHLTA,B195,216;%%

%\cite{Luscher:1994xx}%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\bibitem{Luscher:1994xx}
M.~Luscher,
%``A New approach to the problem of dynamical quarks in numerical simulations of lattice QCD,''
Nucl.\ Phys.\ B {\bf 418} (1994) 637.
%%CITATION = HEP-LAT 9311007;%%


%\cite{Lutest}%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\bibitem{Lutest}
T.~Takaishi and P.~de Forcrand,
%``Simulation of n(f) = 3 QCD by hybrid Monte Carlo,''
Nucl.\ Phys.\ Proc.\ Suppl.\ {\bf 94} (2001) 818.
%%CITATION = HEP-LAT 0011003;%%
R.~Frezzotti and K.~Jansen,
%``A polynomial hybrid Monte Carlo algorithm,''
Phys.\ Lett.\ B {\bf 402} (1997) 328.
%%CITATION = HEP-LAT 9702016;%%
R.~Frezzotti, M.~Hasenbusch, U.~Wolff, J.~Heitger and K.~Jansen,
%``Comparative benchmarks of full QCD algorithms,''
hep-lat/0009027.
%%CITATION = HEP-LAT 0009027;%%
C.~Alexandrou, P.~de Forcrand, M.~D'Elia and H.~Panagopoulos,
%``Efficiency of the UV-filtered multiboson algorithm,''
Phys.\ Rev.\ D {\bf 61} (2000) 074503.
%%CITATION = HEP-LAT 9906029;%%
P.~de Forcrand,
%``Fermionic Monte Carlo algorithms for lattice QCD,''
hep-lat/9702009.
%%CITATION = HEP-LAT 9702009;%%
C.~Alexandrou, A.~Borici, A.~Feo, P.~de Forcrand, A.~Galli, F.~Jegerlehner
and T.~Takaishi,
%``One-flavour QCD at finite temperature,''
Nucl.\ Phys.\ Proc.\ Suppl.\ {\bf 63} (1998) 406.
%%CITATION = HEP-LAT 9710002;%%
I.~Montvay,
%``Multi-bosonic algorithms for dynamical fermion simulations,''
hep-lat/9903029.
%%CITATION = HEP-LAT 9903029;%%
A.~Galli and P.~de Forcrand,
%``Exact local bosonic algorithm for dynamical quarks,''
Nucl.\ Phys.\ Proc.\ Suppl.\ {\bf 53} (1997) 956.
%%CITATION = HEP-LAT 9609013;%%
K.~Jansen,
%``Recent developments in fermion simulation algorithms,''
Nucl.\ Phys.\ Proc.\ Suppl.\ {\bf 53} (1997) 127.
%%CITATION = HEP-LAT 9607051;%%
A.~Borici,
%``Theoretical analysis of multi-boson algorithm with local and global update of bosonic fields,''
hep-lat/9602018.
%%CITATION = HEP-LAT 9602018;%%
A.~Borrelli, P.~de Forcrand and A.~Galli,
%``Non-hermitian Exact Local Bosonic Algorithm for Dynamical Quarks,''
Nucl.\ Phys.\ B {\bf 477} (1996) 809.
%%CITATION = HEP-LAT 9602016;%%
B.~Jegerlehner,
%``Improvements of L\"uscher's local bosonic fermion algorithm,''
Nucl.\ Phys.\ B {\bf 465} (1996) 487;
%%CITATION = HEP-LAT 9512001;%%
%``Study of a new simulation algorithm for dynamical quarks on the APE-100 parallel computer,''
Nucl.\ Phys.\ Proc.\ Suppl.\ {\bf 42} (1995) 879.
%%CITATION = HEP-LAT 9411065;%%
A.~Borici and P.~de Forcrand,
%``Variants of L\"uscher's fermion algorithm,''
Nucl.\ Phys.\ Proc.\ Suppl.\ {\bf 47} (1996) 800;
%%CITATION = HEP-LAT 9509080;%%
%``Systematic errors of Luscher's fermion method and its extensions,''
Nucl.\ Phys.\ B {\bf 454} (1995) 645.
%%CITATION = HEP-LAT 9505021;%%
C.~Alexandrou, A.~Borrelli, P.~de Forcrand, A.~Galli and F.~Jegerlehner,
%``Full QCD with the Luscher local bosonic action,''
Nucl.\ Phys.\ B {\bf 456} (1995) 296.
%%CITATION = HEP-LAT 9506001;%%
B.~Bunk, K.~Jansen, B.~Jegerlehner, M.~Luscher, H.~Simma and R.~Sommer,
%``A New simulation algorithm for lattice QCD with dynamical quarks,''
Nucl.\ Phys.\ Proc.\ Suppl.\ {\bf 42} (1995) 49.
%%CITATION = HEP-LAT 9411016;%%


%\cite{Slavnov}%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\bibitem{Slavnov}
A.~A.~Slavnov,
%``Bosonized formulation of lattice {QCD},''
hep-th/9611154.
%%CITATION = HEP-TH 9611154;%%

%\cite{SLtest}%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\bibitem{SLtest}
T.~Bakeyev and A.~Slavnov,
%``A bosonization algorithm for lattice computer simulations in QCD with dynamical fermions,''
Proc.\ Steklov Inst.\ Math.\ {\bf 228} (2000) 245.
%%CITATION = PSMMB,228,245;%%
T.~Bakeyev,
%``A test of a new simulation algorithm for dynamical quarks,''
JHEP{\bf 9907} (1999) 003.
%%CITATION = HEP-LAT 9905017;%%
T.~Bakeyev, M.~I.~Polikarpov, A.~Slavnov and A.~I.~Veselov,
%``A toy model test of a new algorithm for bosonization of fermion  determinants,''
Theor.\ Math.\ Phys.\ {\bf 113} (1997) 58.
%%CITATION = HEP-LAT 9702002;%%

%\cite{polymer}%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\bibitem{polymer}
C.~Gattringer,
%``The lattice Schwinger model as a discrete sum of filled Wilson loops,''
Nucl.\ Phys.\ B {\bf 559} (1999) 539.
%%CITATION = HEP-LAT 9903021;%%
I.~Montvay,
%``Simulation Of Staggered Fermions By Polymer Averaging,''
Phys.\ Lett.\ B {\bf 227} (1989) 260.
%%CITATION = PHLTA,B227,260;%%
M.~Karowski, R.~Schrader and H.~J.~Thun,
%``Monte Carlo Simulations For Quantum Field Theories Involving Fermions,''
Commun.\ Math.\ Phys.\ {\bf 97} (1985) 5.
%%CITATION = CMPHA,97,5;%%


%\cite{OCCUP}%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\bibitem{OCCUP}
J.~E.~Hirsch, D.~J.~Scalapino, R.~L.~Sugar, R.~Blankenbecler,
Phys.\ Rev. \ B {\bf 26} (1982) 5033.
%%CITATION = PHRVA,B26,5033
I.~Montvay,
%``Bosonized Dynamical Fermions In (3+1)-Dimensions,''
Phys.\ Lett.\ B {\bf 216} (1989) 375.
%%CITATION = PHLTA,B216,375;%%
U.~J.~Wiese,
%``Bosonization and cluster updating of lattice fermions,''
Phys.\ Lett.\ B {\bf 311} (1993) 235.
%%CITATION = HEP-LAT 9210019;%%
S.~Chandrasekharan and U.~Wiese,
%``Meron-cluster solution of a fermion sign problem,''
Phys.\ Rev.\ Lett.\ {\bf 83} (1999) 3116
%%CITATION = COND-MAT 9902128;%%

%\cite{CR2}%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\bibitem{CR2}
M.~Creutz,
%``Evaluating Grassmann integrals,''
Phys.\ Rev.\ Lett.\ {\bf 81} (1998) 3555,
%%CITATION = HEP-LAT 9806037;%%
Phys.\ Rev.\ {\bf B45} (1992) 4650
%%NO ENTRY IN DATABASE

%\cite{Infrared}%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\bibitem{Infrared}
P.~de Forcrand,
%``UV-filtered fermionic Monte Carlo,''
Nucl.\ Phys.\ Proc.\ Suppl.\  {\bf 73} (1999) 822,
%%CITATION = HEP-LAT 9809145;%%
A.~Duncan, E.~Eichten and H.~Thacker,
%``An efficient algorithm for QCD with light dynamical quarks,''
Phys.\ Rev.\ D {\bf 59} (1999) 014505.
%%CITATION = HEP-LAT 9806020;%%
I.~Horvath,
%``Stochastic split determinant algorithms,''
Nucl.\ Phys.\ Proc.\ Suppl.\ {\bf 83} (2000) 804.
%%CITATION = HEP-LAT 9909044;%%

%\cite{NMC}%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\bibitem{NMC}
T.~Bakeyev and P.~de Forcrand,
%``Noisy Monte Carlo revisited,''
Phys.\ Rev.\ D {\bf 63} (2001) 054505.
%%CITATION = HEP-LAT 0008006;%%
%``Noisy Monte Carlo algorithm,''
Nucl.\ Phys.\ Proc.\ Suppl.\ {\bf 94} (2001) 801.
%%CITATION = HEP-LAT 0010006;%%

%\cite{Creutz}%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\bibitem{Creutz}
M.~Creutz,
%``Lattice gauge theory: A retrospective,''
Nucl.\ Phys.\ Proc.\ Suppl.\ {\bf 94} (2001) 219
%%CITATION = HEP-LAT 0010047;%%


\end{thebibliography}

\documentstyle[twoside,fleqn,espcrc2]{article}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\bea}{\begin{eqnarray}}
\newcommand{\eea}{\end{eqnarray}}
\newcommand{\bc}{\begin{center}}
\newcommand{\ec}{\end{center}}

\def\ut{\tilde{U}}
\def\mt{\tilde{M}}

\title{\bf A new method for Monte Carlo simulation of theories
with Grassmann variables.}
\author{T. Bakeyev
\address{Joint Institute for Nuclear Research, 141980 Dubna, Russia}
       }

\begin{document}

\begin{abstract}

A new algorithm for simulation of theories with dynamical fermions is
presented.
The algorithm is based on obtaining the new configuration $\ut$
from the old one $U$ by solving the equation $ M(\ut) \eta= \omega M(U) \eta$,
where $M$ is fermionic operator,
$\eta$ is random Gaussian vector, and $\omega$ is random real number
close to unity.
This algorithm can be used for acceleration of current
simulations in theories with fermions.
\end{abstract}

\maketitle


From the earliest days of Monte Carlo simulations
fermionic fields have caused annoying difficulty,
which stems from their being anticommuting variables.
Although for the most actions in use one can eliminate
fermions by an analytic integration,
the resulting expressions involve determinants
of large matrices, making the numerical simulations expensive.

Over the years many tricks have been developed to solve this
problem. Hybrid Monte Carlo (HMC) \cite{HMC}
is now considered to be the standard simulation tool.
In recent years it
was challenged by the Multiboson (MB) method.
The most popular version of MB,
proposed by M.~L\"uscher \cite{Luscher:1994xx}, has been intensively studied by
many authors (see \cite{Lutest} and references therein).
The MB algorithm, proposed by
A.~Slavnov \cite{Slavnov}, is less known, and was tested in \cite{SLtest}.
Besides HMC and MB, let me mention
some interesting ideas, based on
the polymer \cite{polymer} and Jordan-Wigner occupation number \cite{OCCUP}
representations,
the direct evaluation of Grassmann integrals \cite{CR2},
the separation of low and high eigenmodes of Dirac
operator \cite{Infrared},
and
the direct simulation of loop expansion by means of using the stochastic
estimators \cite{NMC}.

Despite the variety of different approaches to the problem,
there is a common impression that all existing algorithms remain
inefficient \cite{Creutz}. This is clearly seen if we compare the
computational cost for the theories with dynamical fermions from the one side,
and purely bosonic theories from the other side.
Such a situation suggests that one should not stop the attempts to obtain
relatively cheap fermion algorithm.

In this contribution I present a new computational strategy for
treating dynamical fermions in Monte Carlo simulations.
It has the virtues of exactness, nonlocality and finite step-size of update
Unlike the other exact algorithms,
the CPU-time required for the update in this algorithm grows
only linearly with the volume $V$ of the system.

Suppose that we aim at sampling the partition function
of theory with two flavors of degenerate fermion fields:
\be {\cal Z}_{ferm} = \int Det \;\Bigl( M^\dagger [U] M[U] \Bigr)
\; dU \label{par_func}\ee
where $M$ is fermion operator, $U$ denotes the bosonic fields
coupled to fermions.
(For simplicity of formulae I do not include the purely
bosonic contribution.)
Then starting from old configuration $U$, one executes the following
instructions:

\begin{itemize}
\item generate random vector $\eta$ with Gaussian
distribution:
$ \;\;\; P_G[\eta] = Z^{-1}_G e^{-|\eta|^2} $
\item generate
$\;\omega \in \Bigl[1-\epsilon\; ;\; \frac{1}{1-\epsilon}\Bigr]\;$
with the probability:
${\cal P}[\omega,\eta]\propto min\; \bigl(1/\omega^2,
e^{\bigl(1 - \frac{1}{\omega^2}\bigr)|\eta|^2}\bigr)$
\item find the new configuration $\ut$ by solving the equation:
\be \mt \eta = \omega M\eta   \label{basic}\ee
\end{itemize}
where $0 \le \epsilon < 1$ is an algorithmic parameter.
Here and below I use the following short notations:
$\mt \equiv M[\ut] ;\; M \equiv M[U]\;$.

In Ref. \cite{Bakeyev:2001wb} it was proved that as far as the solution
of eq.(\ref{basic}) is found in a reversible way
\footnote{i.e. the probability to obtain
configuration $\ut$ starting from $U$ at any $\omega$ should be equal
to probability to obtain $U$ starting from $\ut$ at $1/\omega$},
the algorithm satisfies the detailed balance condition.

It is clear that choosing the procedure of solving
the eq.(\ref{basic}) is crucial in this approach. Firstly, because
obtaining the solution $\ut$ is the only
computationally expensive ingredient of the algorithm.
Secondly, because not every procedure of solving
the eq.(\ref{basic}) can provide ergodicity.
Finding a good procedure of solving the eq.(\ref{basic}) presents
a serious challenge for the future investigations. The
potential gain worth the efforts: e.g. finding
some analytic solutions can give us a very cheap fermionic algorithm,
comparable in cost with the algorithms for purely bosonic theories.

Making the first attempt to test the new algorithm, I solved
the eq.(\ref{basic}) numerically
by using the local iterative minimization of the quantity
\be R\equiv  |\; (\mt-\omega M)\eta \; |^2 \label{r_min}\ee
for fixed $U,\omega,\eta$. Starting from $\ut = U$,
the minimization proceeds until $R<\delta$ is reached, where $\delta$
determines the accuracy of solution.

The tests were done for SU(3) QCD at $8^4$ lattice for the
partition function (\ref{par_func}) with
\be M[U]=1 - k^2 D_{eo}D_{oe}\ee
where $\; D[U]$ is Wilson difference operator.
The even-odd preconditioning was used to reduce the computational cost.
The hopping parameter was $k=0.2$, which gave the plaquette
${\cal P} = 0.0089(1)$.
The minimization of functional (\ref{r_min}) was implemented by making
the random moves in each color direction for all links lexicographically.

I checked the existence of solutions of eq.(\ref{basic}) at
$\omega \in [0.9, 1.1]$ for typical equilibrated
configurations $U$. It was found that for considered
$\omega$ one always finds some solution.
On the average solving the eq.(\ref{basic}) at $\epsilon = 0.05$
with precision
$\delta = 5*10^{-7}$ required 42 minimization iterations
(it was checked, that improving the precision did not affect the results).
Since the cost of one iteration is roughly equal to $2*N_{generator}=16$
multiplications
by matrix $M$, the total cost of finding the solution
was approximately 670 matrix multiplication.
This has to be compared with the cost for generating one trajectory with HMC.
At step-size $\tau = 0.033$ and trajectory
length equal to $1$ (this ensures 70\% acceptance),
one trajectory cost approximately
4420 matrix multiplications, i.e. $\approx 6.6$ times more expensive than
solving eq.(\ref{basic}).

My conclusion was that the usage of minimization
of functional (\ref{r_min}) for solving the eq.(\ref{basic})
is not favorable, because the algorithm became nonergodic.
The problem is that probability
${\cal P}[\omega,\eta]$ to accept $\omega<1$ is
strongly suppressed by large factor in the exponent, if $\omega$
is far from 1.
Indeed, the squared norm of Gaussian vector can be estimated as
$\; |\eta|^2 \sim I_{dof}$, where $ I_{dof} $ is the dimensionality
of vector space on which $\eta$ is defined.
For even-odd preconditioned SU(3) QCD at $8^4$ lattice one has
$I_{dof} = 24576$!
Therefore, using the new algorithm alone,
one is restricted to choose between two unfavorable possibilities:
1) Using $\epsilon \sim 1/I_{dof} $.
It was observed that at such small values of $\epsilon$
the evolution of $U$ fields becomes
very slow, because the new configuration $\ut$ always lies too close to
the old one;
2) Using $\epsilon \gg 1/I_{dof}$. Then $\omega < 1$ is almost never
accepted, and the configurations are becoming smother.

In the second case the trouble appears due to the nonergodicity of algorithm
at large $\epsilon$. Nevertheless, even at large $\epsilon$
our algorithm can be used for acceleration of other fermionic
algorithms like HMC or MB, since the ergodicity is provided
by them.

I tested the combination of the new algorithm and HMC. After each
HMC trajectory the global move in configuration space was
implemented
by using our algorithm at $\epsilon = 0.05$.
The efficiency of this algorithmic
mixture was estimated by measuring
the autocorrelation times for the plaquette. The results were:
1) Pure HMC, 1000 trajectories: $\tau_{int}=1.4(2)$;
2) HMC+new alg., 1000 (traj + $\omega$-update): $\tau_{int}=0.7(1)$.
Note, that autocorellation time was reduced almost at no cost,
because the global update of the new algorithm is much cheaper, than HMC
trajectory.
Of course, these tests are not very illustrative, since the
autocorrelation times for theory with pure fermionic term at this kappa
value are rather small.
The further tests in models of practical relevance are needed.

Let me emphasize, that nonergodicity of the new algorithm
was produced by the particular way of solving the eq.(\ref{basic}),
which forces us to use large $\epsilon$.
One should think of finding some other procedure,
which allows to use $\epsilon = 0$ (or $\epsilon \approx 0$).
Unfortunately, in this case the search for the solution of eq.(\ref{basic})
by minimizing the functional (\ref{r_min}) does not work,
because one immediately gets the trivial solution $\ut = U$.
However, in typical case (e.g. SU(3) QCD), the solutions of eq.(\ref{basic})
are highly degenerate
\footnote{
In SU(3) QCD the expression (\ref{basic}) for even-odd preconditioned
matrix provides $12V$ real equations for $32V$ variables.},
and the subspace of nontrivial solutions is not empty.
Finding the reversible procedure
of nontrivial solving the equation $\mt\eta = M\eta$,
which samples the configuration space fast enough,
is the perspective subject for future investigations.

I conclude that the efficiency of the new fermion algorithm 
drastically 
depends on the clever choice of procedure of solving the 
eq.(\ref{basic}). Finding the new efficient way of solving this 
equation can provide a very cheap simulation method.

{\bf Acknowledgments:}
I am very grateful to the Organizers of LATTICE2001
for generous help and support.
I thank Ph. de Forcrand for helpful comments.
This work was supported by grant for young scientists
\# 01-01-06089, Russian Basic Research Fund
\# 99-01-00190 and president grant for support of
scientific schools \# 0015-96046.

\end{document}